3. A neutral mutation occurs in a monomorphic population of size 5000, so that in the next generation there are 4999 wild-type individuals and one mutant. The efective population size of this population is 1000. Assuming that no other mutations occur in the near future, what is the probability that this new neutral mutant will fix in the population?
Answer #1

Effective population size (Ne) = 1000

Probability of new neutral mutant to fix in the population, p = 1/2Ne = 1/2 x1000 = 1/2000.

New mutant fixation in a population probability = 1/2000 = 0.0005

probability in percentage = 0.0005 x 100 = 0.05 %